the width of a rectangle is 20 cm more than the length. The perimeter is 240. Find the length and width.
Answer by PeanutTheCat(5)

Write Let Statements
- Let L = length
- Let 20 + L = width

Step 1: Plug in to Perimeter "Formula"
P = 2(L) + 2(w)
240 = 2L + 2(20 + L)
240 = 2L + 40 + 2L
240 = 4L + 40
200 = 4L
L = 50 cm

Step 2: Find Width
Width = 20 + L
Width = 20 + 50
Width = 70

Final Answers:
- Length = 50 cm
- Width = 70 cm

To Check:
P = 2L + 2w
240 = 2(50) + 2(70)
240 = 240
Width (70 cm) is 20 cm more than length (50 cm)